OpenStudy (anonymous):

A hard candy has a radius of 9 mm. The candy has a spherical soft center with a radius of 6 mm, with a hard shell surrounding it. What is the volume of the hard shell? Use 3.14 to approximate pi and round the answer to two decimal places.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@saynabahmed13

OpenStudy (anonymous):

Hi i can help

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

V = (4/3)π(r)^3 V = (4/3)3.14(9-6)^3 V = (4/3)3.14(3)^3 V = 4/3(9.42)^3 V = 12.56^3 V = 1981.385216 Answer: 1981.39 mm3

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

I think I will recheak

OpenStudy (anonymous):

I think it is the right answer give it a try

OpenStudy (anonymous):

thanks bro

OpenStudy (anonymous):

do you play billards

OpenStudy (anonymous):

I am a girl lol

OpenStudy (anonymous):

But we can be friends :)

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

its ok can u fan me

OpenStudy (anonymous):

I mean medal

OpenStudy (anonymous):

im new how do i do that?

OpenStudy (anonymous):

u just did it lol

OpenStudy (anonymous):

oh. haha

OpenStudy (anonymous):

we can be friends on open study

OpenStudy (anonymous):

anyways P.S I hate to learn

OpenStudy (anonymous):

ok. thats cool. Do you play billards because your profile has pool balls on it

OpenStudy (anonymous):

no Its just a pic on my profile